# Data-driven rainfall prediction at a regional scale: a case study with Ghana

### v2

@@ 4 Model evaluation
We evaluate and compare four models: $\mathsf{UNET}_{12}$, $\mathsf{UNET}_{30}$, the 18h lead-time predictions of the ECMWF obtained from TIGGE, and denoted as $\mathsf{NWP}$, and model $\mathsf{Ens}$ that is the weighted average of $\mathsf{UNET}_{12}$ and $\mathsf{NWP}$. We use the climatology prediction as reference, and denote it as $\mathsf{CLIM}$. The climatology rainfall prediction at location $i$ and date $t$ is defined as the average of all the 24h rainfall values in the GPM-IMERG data set occurring at location $i$, and on same day of the year as $t$ (plus two days prior and two days after). We compare these models using the mean absolute error (MAE), as well as the precision, recall and F1-score for rainfall event detection.
4.1 Prediction error
For a given model $m\in\{\mathsf{CLIM},\mathsf{UNET}_{12},\mathsf{UNET}_{30},\mathsf{NWP},\mathsf{Ens}\}$, we write $\hat{y}_{it^{\prime}}(m)$ to denote the predicted rainfall at location $i\in 64\times 64$, and at time $t^{\prime}$ by model $m$. The mean absolute prediction error (MAE) of model $m$ at pixel $i$ is defined as
@@ 4.1 Prediction error
where the average is taken over the test dataset $\mathcal{D}^{{}^{\prime\prime}}$. This metric can also be viewed as a special case of the CRPS (Gneiting and Raftery (2007)) with one forecast value. Using the NWP prediction as reference, we compute the pixel-by-pixel prediction error skill of model $m$ as
$\mathsf{Skill}(i,m)=\frac{\ell(i,\mathsf{NWP})-\ell(i,m)}{\ell(i,\mathsf{NWP})},$ (5)
which offers a visual evaluation of model $m$. As defined, $\mathsf{Skill}(i,m)$ is positive if model $m$ performs better than the ECMWF forcast at $i$, and $\mathsf{Skill}(i,m)$ is nonpositive otherwise.
@@ 4.2 Precision, recall, and F1-score
We also compare the models in their ability to correctly predict rainfall events. Given a threshold level $\tau$, and a model $m\in\{\mathsf{CLIM},\mathsf{UNET}_{12},\mathsf{UNET}_{30},\mathsf{NWP},\mathsf{Ens}\}$, its precision and recall at pixel $i$ are defined respectively as
@@
We further combine the precision and the recall into a $(\mathcal{F}_{1})$ score
@@
We compute $\mathcal{R}(m)$ and $(\mathcal{F}_{1})(m)$ similarly. We compute these statistics with two threshold values: $\tau=0.5$mm which allows us to compare the models in their ability to predict rainfall events, and $\tau=10$mm to evaluate the ability of the models to prediction heavy rainfall events (a larger value of $\tau$ would lead to rare-events and high-variance estimates).
@@ 4.3 Probing the importance of the input variables
One of the main limitations of deep learning models is their lack of interpretability. With the development and adoption of these techniques in Earth science research, there is a growing need for interpretable/explanable deep learning models (Bommer et al., 2024). We focus here on our model $\mathsf{UNET}_{12}$, and we propose a novel methodology to evaluate the importance of its input predictors. Let $\mathcal{F}_{\hat{W}}$ denote the fitted model $\mathsf{UNET}_{12}$. We recall that the input feature ${\bf x}$ is of dimension ${\bf x}\in\mathbb{R}^{K\times I}$, where $K$ denote the number of variables, and $I$ the number of pixels (we have collapsed the longitude and latitude dimensions into one pixel dimension). One common approach that one could use to measure the importance of a given feature input in this setting is the sensitivity of the predicted rainfall with respect to that input (Baehrens et al., 2010; Bommer et al., 2024). Specifically, the sensitivity of the predicted rainfall to the input variable $k$ is defined as
@@
where the average is over the test dataset $\mathcal{D}^{{}^{\prime\prime}}$. (Bommer et al., 2024) presented several variations of this metric that have been used in the literature. We note however that these sensitivity metrics are defined without any reference to model performance. Hence a variable could have high sensitivity without being essential for accurate prediction.
To address this issue, we propose a novel methodology built on the idea of posterior predictive checking, a powerful tool in Bayesian data analysis for model checking and validation (Gelman et al. (2004)). The basic idea is the use of the posterior predictive distribution as a data generating process for the test data. Based on that data generating process we then perform a variable selection to determine which input variable best explains the observed performance on the test data. To be more specific, suppose for a moment that we have fitted our deep neural network model in the Bayesian framework with a posterior distribution $\Pi(\cdot|\mathcal{D}^{\prime})$. Then the posterior predictive distribution of a data point $\mathbf{x}_{t-h}$ in the test set $\mathcal{D}^{{}^{\prime\prime}}$ would be defined as
@@
which gives the relative importance of each variable $\delta_{j}$ in explaining the performance of our fitted model $\Pi(W|\mathcal{D}^{\prime})$ on the test data $\mathcal{D}^{{}^{\prime\prime}}$. We note that the posterior predictive distribution $\Pi(\delta|\mathcal{D}^{\prime}\cup\mathcal{D}^{{}^{\prime\prime}})$ in (6) depends both on the test set $\mathcal{D}{{}^{\prime\prime}}$ and the training set $\mathcal{D}^{\prime}$. We note also that the test set $\mathcal{D}^{{}^{\prime\prime}}$ is typically not very large, therefore we do not expect $\Pi(\delta|\mathcal{D}^{\prime}\cup\mathcal{D}^{{}^{\prime\prime}})$ to concentrate, in the sense of putting high probability on a small number of variables. Instead it provides a ranking of the relative importance of the input variables in explaining the performance of the fitted model on the test data.
@@
$\textsf{L}(\delta;\mathcal{D})=\sum_{t\in\mathcal{D}^{{}^{\prime\prime}}}\left\|\textbf{y}_{t}-\mathcal{F}_{\widehat{W}}(\delta\odot\mathbf{x}_{t-h})\right\|_{1}.$
We note that L depends both on the test set $\mathcal{D}{{}^{\prime\prime}}$ and the training set $\mathcal{D}^{\prime}$. We sample from (7) using a straightforward Gibbs sampling. For an introduction to MCMC and the Gibbs sampler, we refer the reader to (Brooks et al., 2011). Specifically, given $\delta\in\{0,1\}^{K}$, and $1\leq j\leq K$, let $\delta^{(j,1)}$ (resp. $\delta^{(j,0)}$) be the element of $\{0,1\}^{K}$, that is equal to $\delta$ except perhaps on component $j$, where $(\delta^{(j,1)})_{j}=1$ (resp. $(\delta^{(j,0)})_{j}=0$). Then under the joint (7), the conditional distribution of $\delta_{j}$ given the remaining components denoted $\delta_{-j}$ and given $\mathcal{D}$ is a Bernoulli distribution $\textbf{Ber}(q_{j})$, with probability $q_{j}$ given by
@@ 5 Results and Discussion
We evaluate the performance of the models using the metrics presented above. As a start, Figure 3 shows some prediction samples from the models.
@@
Figure 3: Sample forecasts ($\mathsf{NWP}$, $\mathsf{CLIM}$, $\mathsf{UNET}_{30}$, $\mathsf{UNET}_{12}$, and $\mathsf{Ens}$), along with the corresponding ground truth (GT) from GPM-IMERG.
5.1 Comparison in terms of mean absolute error
@@
Table 2 shows the overall mean and standard deviation of the mean absolute errors computed across all pixels in the test dataset. The results show that in terms of the mean absolute error, the numerical weather prediction model $\mathsf{NWP}$ model is comparable to the climatology $\mathsf{CLIM}$, but $\mathsf{UNET_{12}}$ and the ensemble model show clear improvement. The fact that the numerical weather prediction model does not outperform the climatology forecast under the tropics is well-documented as we discussed in the introduction (Vogel et al. (2020)). However, climatology forecasts are broad averages and therefore lack precision as we will see below.
Figure 4 shows the MAE skill maps of the models compared to $\mathsf{NWP}$ as given in (5). White areas holds negative values and indicate locations at which the NWP has better MAE. The maps show the NWP significantly under-performs $\mathsf{UNET_{12}}$, Ens, (and $\mathsf{UNET_{30}}$ to a lesser extent), particularly in the coastal south and the north.
@@
Table 2: Means and standard deviations of the Mean absolute errors
CLIM NWP $\mathsf{UNET}_{30}$ $\mathsf{UNET}_{12}$ Ens
Mean 3.89 3.92 3.83 3.74 3.63
Std 1.15 1.12 1.28 1.13 1.10
Figure 4: MAE skill maps. We treat the spatial locations as independent. Each skill value is computed as in (5)
5.2 Comparing the biases of $\mathsf{NWP}$ and $\mathsf{UNET}_{12}$ predictions
@@
Here we further investigate the structure of the prediction biases of the models $\mathsf{NWP}$ and $\mathsf{UNET}_{12}$. We combine all the pixels of all the test samples together, and look at the proportion of pixels with bias (defined as the difference between the predicted rainfall and the corresponding GPM-IMERG estimate) falling within each of the following ranges $(-0.1,0)$, $(0,0.1)$, $(-1,0)$, and $(0,-1)$. The results are presented in Table 3. We note that $\mathsf{UNET}_{12}$ has 50% more bias values in the interval $(-0.1,0.1)$ than $\mathsf{NWP}$, and 22% more bias values in the interval $(-1,1)$. These statistics support again our conclusion that the DL model $\mathsf{UNET}_{12}$ performs better than $\mathsf{NWP}$. We note however that $\mathsf{UNET}_{12}$ has a strong tendency to underestimate the rainfall values, whereas $\mathsf{NWP}$ has a tendency to overestimate the rainfall values. These observed bias patterns has motivated the introduction of our ensemble forecast model Ens that combines $\mathsf{UNET}_{12}$ and $\mathsf{NWP}$.
Table 3: Proportion of pixels (in percentage) in each range
Model Proportions
(-0.1, 0) (0, 0.1) (-1, 0) (0, 1)
$\mathsf{NWP}$ 11.99 16.72 16.39 32.74
$\mathsf{UNET}_{12}$ 34.07 10.00 42.26 17.69
@@ 5.3 Comparison in terms of precision and recall
Correctly matching the tail events of the rainfall distribution has significant practical values. Here we compare our forecast models in their ability to correctly predict rainy days and heavy rainy days (days with 24h rainfall larger than $0.5$mm, and $10$mm respectively). Using the GPM-IMERG as ground truth, we compare the models in terms of precision (P), recall (R) and F1 score as described in Section 4. Table 4 summarizes the overall performance of each model. We first consider the detection of rainy days (at threshold $\tau=0.5$mm). As expected, due to broad averaging, the CLIM forecast suffers from low precision ($0.45$), as it frequently misclassifies non-rain events as rain. Similarly, the NWP model also shows a strong recall ($88\%$) but a relatively low precision ($0.52$). This suggests that the NWP model captures much of the rainfall signal, but also has some amount of spurious phenomena driving its forecasts. In contrast, the data-driven models have lower recall, which suggests that these models are missing some of the rainfall signal (the physics), however they have better precision. The ensemble model (Ens) is able to leverage the complementary strengths of the two approaches and scores better. The location-wise distribution of P and R across the area of interest can be further visualized in Figures 5 and 6, respectively.
@@
Precision Recall F1 Precision Recall F1
$\mathsf{CLIM}$ 0.45 0.97 0.61 0.17 0.02 0.04
$\mathsf{NWP}$ 0.52 0.88 0.65 0.27 0.25 0.26
$\mathsf{UNET}_{30}$ 0.59 0.57 0.58 0.23 0.14 0.17
$\mathsf{UNET}_{12}$ 0.62 0.64 0.63 0.31 0.26 0.28
$\mathsf{Ens}$ 0.61 0.79 0.69 0.32 0.25 0.28
Figure 5: Locationwise precision (P) of the models at threshold $\tau=0.5$mm
Figure 6: Locationwise recall (R) of the models at threshold $\tau=0.5$mm
5.3.1 Detection of heavy rainfall
We also compare the models through their predictions of heavy rainy days, particularly on how such predictions agree with GPM-IMERG. We set the threshold for heavy rain to $\tau=10$mm. The performance of the models is much worst. This suggests that more work is needed to better predict large and extreme rainfall events. Figures 7 and 8 show the location-wise distribution of P and R across the area of interest (AOI). From these figures we see again that $\mathsf{UNET_{12}}$ (and $\mathsf{UNET_{30}}$ to a lesser extent) performs better than NWP.
Figure 7: Locationwise precision (P) of the models at threshold $\tau=10$mm
Figure 8: Locationwise recall (R) of the models at threshold $\tau=10$mm

### v3

@@ 4 Model evaluation
We evaluate and compare four models: UNET12, UNET18, the 18-hour lead-time predictions of ECMWF, obtained from TIGGE and denoted as NWP, and the hybrid model HYB, which is a combined version of UNET18 and NWP. We use the climatology prediction as reference, and denote it as CLIM. The climatology-based rainfall prediction at location $i$ and date $t$ is defined as the average of all the 24h rainfall values from the GPM-IMERG dataset recorded in previous years at the same location $i$, considering the same calendar day $t$, as well as two days before and two days after. We also use this ensemble of past rainfall values as an empirical CDF for the purpose of uncertainty quantification. We compare these models using the mean absolute error (MAE), the CRPS as defined above, as well as the precision, recall, and F1-score for rainfall event detection.
4.1 Prediction error in terms of MAE and CRPS
For a given model $m\in\{\rm{CLIM,UNET_{12},UNET_{18},NWP,HYB}\}$, we write $\hat{y}_{it^{\prime}}(m)$ to denote the predicted rainfall at location $i\in 1\dv 64\times 1\dv 64$, and at time $t^{\prime}$ by model $m$. The mean absolute prediction error (MAE) of model $m$ at pixel $i$ is defined as
@@ 4.1 Prediction error in terms of MAE and CRPS
where the average is taken over the test dataset $\mathcal{D}^{{}^{\prime\prime}}$. Similarly, let $\hat{F}_{i,t^{\prime}}^{(m)}(\cdot)$ denote the estimated cdf for method $m$ at time $t^{\prime}$ and pixel $i$. The CRPS of model $m$ at pixel $i$ is defined as:
${\rm CRPS}(i,m)=\frac{1}{|\mathcal{D}^{{}^{\prime\prime}}|}\sum_{t^{\prime}\in\mathcal{D}^{{}^{\prime\prime}}}{\rm CRPS}(\hat{F}_{i,t^{\prime}}^{(m)},y_{i,t^{\prime}}).$
In our setting, all the cdfs $\hat{F}_{i,t^{\prime}}^{(m)}(\cdot)$ are discrete, and the CRPS can be efficiently computed as in (Jordan et al., 2019). The CRPS reduces to the MAE for ensembles with a single element. Using the CLIM prediction as a reference, we compute the pixel-by-pixel CRPS error skill of model $m$ as
$\rm{Skill}(i,m)=\frac{{\rm CRPS}(i,{\rm CLIM})-{\rm CRPS}(i,m)}{{\rm CRPS}(i,{\rm CLIM})},$ (6)
which offers a visual evaluation of model $m$. As defined, $\rm{Skill}(i,m)$ is positive if model $m$ performs better than the CLIM forecast at $i$, and $\rm{Skill}(i,m)$ is nonpositive otherwise.
@@ 4.2 Precision, recall, and F1-score
We also compare the models in their ability to correctly predict rainfall events. Given a threshold level $\tau$, and a model $m\in\{\rm{CLIM,UNET_{12},UNET_{18},NWP,HYB}\}$, its precision and recall at pixel $i$ are defined respectively as
@@
The recall statistic $\mathcal{R}_{i}(m)$ measures the proportion of days in the test set with rain amount larger than $\tau$ correctly detected by method $m$ (at location $i$). The precision $\mathcal{P}_{i}(m)$ is the true positive rate of method $m$ at location $i$: among the days that method $m$ assigned as having rain amount larger than $\tau$, the proportion of them that indeed have rain amount larger than $\tau$. We further combine the precision and the recall into a $(\mathcal{F}_{1})$ score
@@
We compute $\mathcal{R}(m)$ and $(\mathcal{F}_{1})(m)$ similarly. We compute these statistics with two threshold values: $\tau=0.5$ mm, which allows us to compare the models in their ability to predict rainfall events, and $\tau=10$ mm to evaluate the ability of the models to predict heavy rainfall events. For the threshold $\tau=0.5$ mm, 34.9% of all grid cells have rainfall above this threshold, while 46.1% of rainy grid cells (i.e., rainfall greater than 0 mm) exceed it. At $\tau=10$ mm, the corresponding values are 10.0% and 13.2%, respectively. Overall, 75.6% of grid cells in the dataset record nonzero rainfall.
@@ 4.3 Probing the importance of the input variables
One of the main limitations of deep learning models is their lack of interpretability. With the development and adoption of these techniques in Earth science research, there is a growing need for interpretable/explainable deep learning models (Bommer et al., 2024). We focus here on our model $\rm{UNET}_{12}$, and we propose a novel methodology to evaluate the importance of its input predictors. Let $\mathcal{F}_{\hat{W}}$ denote the fitted model $\rm{UNET}_{12}$. We recall that the input feature ${\bf x}$ is of dimension ${\bf x}\in\mathbb{R}^{K\times I}$, where $K$ denote the number of variables, and $I$ the number of pixels (we have collapsed the longitude and latitude dimensions into one pixel dimension). A common approach that one could use to measure the importance of a given input feature in this setting is the sensitivity of the predicted rainfall with respect to that input (Baehrens et al., 2010; Bommer et al., 2024). Specifically, the sensitivity of the predicted rainfall to the input variable $k$ is defined as
@@
where the average is over the test dataset $\mathcal{D}^{{}^{\prime\prime}}$. Bommer et al. (2024) presented several variations of this metric that have been used in the literature. We note however, that these sensitivity metrics are defined without any reference to model performance. Hence, a variable could have high sensitivity without being essential for accurate prediction.
To address this issue, we propose a novel methodology built on the idea of posterior predictive checking, a powerful tool in Bayesian data analysis for model checking and validation (Gelman et al., 2004). The basic idea is the use of the posterior predictive distribution as a data-generating process for the test data. Based on that data-generating process, we then perform a variable selection to determine which input variable best explains the observed performance on the test data. To be more specific, suppose for a moment that we have fitted our deep neural network model in the Bayesian framework with a posterior distribution $\Pi(\cdot|\mathcal{D}^{\prime})$. Then the posterior predictive distribution of a data point $\mathbf{x}_{t-h}$ in the test set $\mathcal{D}^{{}^{\prime\prime}}$ would be defined as
@@
which gives the relative importance of each variable $\delta_{j}$ in explaining the performance of our fitted model $\Pi(W|\mathcal{D}^{\prime})$ on the test data $\mathcal{D}^{{}^{\prime\prime}}$. We note that the posterior predictive distribution $\Pi(\delta|\mathcal{D}^{\prime}\cup\mathcal{D}^{{}^{\prime\prime}})$ in (7) depends both on the test set $\mathcal{D}{{}^{\prime\prime}}$ and the training set $\mathcal{D}^{\prime}$. We note also that the test set $\mathcal{D}^{{}^{\prime\prime}}$ is typically not very large; therefore, we do not expect $\Pi(\delta|\mathcal{D}^{\prime}\cup\mathcal{D}^{{}^{\prime\prime}})$ to concentrate, in the sense of putting high probability on a small number of variables. Instead, it provides a ranking of the relative importance of the input variables in explaining the performance of the fitted model on the test data.
@@
$\mathsf{L}(\delta;\mathcal{D})=\sum_{t\in\mathcal{D}^{{}^{\prime\prime}}}\left\|\textbf{y}_{t}-\mathcal{F}_{\widehat{W}}(\delta\odot\mathbf{x}_{t-h})\right\|_{1}.$
We note that $\mathbf{L}$ depends both on the test set $\mathcal{D}{{}^{\prime\prime}}$ and the training set $\mathcal{D}^{\prime}$. We sample from (8) using a straightforward Gibbs sampler. For an introduction to MCMC and the Gibbs sampler, we refer the reader to Brooks et al. (2011). Specifically, given $\delta\in\{0,1\}^{K}$, and $1\leq j\leq K$, let $\delta^{(j,1)}$ (resp. $\delta^{(j,0)}$) be the element of $\{0,1\}^{K}$, that is equal to $\delta$ except perhaps on component $j$, where $(\delta^{(j,1)})_{j}=1$ (resp. $(\delta^{(j,0)})_{j}=0$). Then under the joint (8), the conditional distribution of $\delta_{j}$ given the remaining components denoted $\delta_{-j}$ and given $\mathcal{D}$ is a Bernoulli distribution $\textbf{Ber}(q_{j})$, with probability $q_{j}$ given by
@@ 5 Results and Discussion
We evaluate the performance of the models using the metrics presented above. As a start, Figure 3 shows some random ground truth rainfall images, and their predictions from the models under comparison.
@@
Figure 3: Sample forecasts ($\rm{CLIM,NWP,UNET_{18},UNET_{12},HYB}$), along with the corresponding ground truth (GT) from GPM-IMERG.
5.1 Comparison in terms of mean absolute error and CRPS
@@
Table 2 shows the overall mean and standard deviation of the MAE and CRPS computed across all pixels in the test dataset. The results show that in terms of the MAE, the $\rm{NWP}$ model is comparable to the climatology $\rm{CLIM}$, but $\rm{UNET_{12}}$, $\rm{UNET_{18}}$, and the HYB model show clear improvement. The fact that the numerical weather prediction model does not outperform the climatology forecast under the tropics is well-documented, as we discussed in the introduction (Vogel et al., 2020). However, climatology forecasts are broad averages and therefore lack precision, as we will see below.
Figure 4 presents the location-wise CRPS maps for all comparative approaches. Overall, both UNET models perform better than $\rm{NWP}$ and $\rm{CLIM}$ across most regions, while the $\rm{NWP}$ model demonstrates stronger performance in the southeastern corner of the area of interest. Similarly, Figure 5 shows the CRPS skill maps of the models compared to $\rm{CLIM}$ as given in (6). White areas hold negative values and indicate locations at which the $\rm{CLIM}$ has better CRPS. The maps are consistent with the summary results on Table 2 and show that the $\rm{CLIM}$ significantly under-performs $\rm{UNET_{12}}$ and $\rm{UNET_{18}}$. The hybrid method HYB has improved MAE as shown on Table 2 but has comparatively poor CRPS score. This is possibly due to our inability to apply EasyUQ to HYB as we explain in section 3-3.1. Nevertheless, we note in Figure 5 that the hybrid model performs better than NWP and UNET12 in the northern and south-east regions, but worse in the central region.
@@
Table 2: Means and standard deviations (SD) of the Mean absolute errors (MAE) and Continuous Ranked Probability Score (CRPS)
Model Means of MAE SD of MAE Means of CRPS SD of CRPS
CLIM 3.90 1.15 2.98 1.27
NWP 3.92 1.00 3.94 1.16
$\rm{UNET_{18}}$ 3.81 1.22 2.71 0.96
$\rm{UNET_{12}}$ 3.74 1.13 2.72 0.96
HYB 3.69 1.01 3.01 0.80
Figure 4: Location-wise CRPS map obtained using $\rm{CLIM,NWP,UNET_{18},UNET_{12},HYB}$ models
Figure 5: CRPS skill maps. We treat the spatial locations as independent. Each skill value is computed as in (6). For increased sharpness of the figures, the skill values have been discretized into three categories: 1 (better than CLIM), 0 (equal to CLIM), and -1 (worse than CLIM). The UNET models show superior skill in most regions, while the NWP forecasts perform better in the southeast corner of the AOI.
5.2 Comparing the biases of NWP and $\rm{UNET_{18}}$
@@
Here we further investigate the structure of the prediction biases of the models NWP and $\rm{UNET_{18}}$. We define the bias as the difference between the predicted rainfall and the corresponding ground truth. Clearly, the closer the bias is to zero, the better. We combine all the pixels of all the test samples together, and look at the proportion of pixels with bias falling within each of the following ranges $(-\infty,-1],(-1,-0.1],(-0.1,0]$, $(0,0.1]$, $(0.1,1]$, and $(1,\infty)$. The results are presented in Table 3. We note that the proportion of bias values produced by $\rm{UNET_{18}}$ that fall in the combined interval $(-0.1,0.1)$ is $70\%$ higher than for NWP. We also note that $\rm{UNET_{18}}$ tends to have more negative bias, compared to NWP which is positively biased. This observation has initially motivated our construction of the hybrid method.
Although the individual pixel-wise forecasts are not independent, one can perform a homogeneity chi-square test (see e.g. Agresti (2007) Section 2.4), to check whether the two models have similar probabilities of landing their forecast biases in these 6 intervals. Let $p_{m,c}$ be the probability that at any given pixel, method $m\in\{{\rm NWP,UNET}\}$ produces a forecast with bias in interval $c\in\{1,\ldots,6\}$, where we use ${\rm UNET}$ as a shorthand for ${\rm UNET_{18}}$. Let $N_{m,c}$ be the number of times method $m\in\{{\rm NWP,UNET}\}$ produces a forecast with bias in interval $c$ as given in Table 3. Then, under the assumption that individual pixel-wise forecasts are independent, and under the null assumption that $p_{{\rm NWP},c}=p_{{\rm UNET},c}$ for $c=1,\ldots 6$, the statistic
$S=\sum_{c=1}^{6}\frac{\left(N_{{\rm UNET},c}-N_{{\rm NWP},c}\right)^{2}}{\left(N_{{\rm UNET},c}+N_{{\rm NWP},c}\right)},$
follows asymptotically a chi-square distribution with $5$ degrees of freedom. The value of $S$ is $S=39,426.78$, with p-value $=0.0$ (up to the machine epsilon). The most salient difference between the two distributions is that ${\rm UNET}_{18}$ assigns higher probability to the combined interval $(-0.1,0.1]$. These results thus support the conclusion that the DL model $\rm{UNET_{18}}$ shows better performance compared to NWP.
Table 3: Counts and Proportions (in percentage) of pixels in each range
Model Counts and Proportions
($\infty$- ∞, -1] (-1, -0.1] (-0.1, 0] (0, 0.1] (0.1, 1] (1, $\infty$)
$\rm{NWP}$ 43237 (19.19) 9546 (4.24) 24404 (10.83) 32850 (14.58) 39088 (17.35) 76155 (33.80)
$\rm{UNET_{18}}$ 47820 (21.23) 18540 (8.23) 17525 (7.78) 79252 (35.18) 19671 (8.73) 42472 (18.85)
@@ 5.3 Comparison in terms of precision and recall
Correctly matching the tail events of the rainfall distribution has significant practical value. Here we compare our forecast models in their ability to correctly predict rainy days and heavy rain days (days with 24h rainfall larger than $0.5$mm, and $10$mm respectively). Using the GPM-IMERG as ground truth, we compare the models in terms of precision (P), recall (R), and F1-score as described in Section 4. Table 4 summarizes the overall performance of each model. We first consider the detection of rainy days (at threshold $\tau=0.5$mm). As expected, due to broad averaging, the CLIM forecast suffers from low precision ($0.45$), as it frequently misclassifies non-rain events as rain. Similarly, the $\rm{NWP}$ model also shows a strong recall ($88\%$) but a relatively low precision ($0.49$). This suggests that the $\rm{NWP}$ model captures much of the rainfall signal, but also has some level of spurious effects driving its forecasts. In contrast, the data-driven models have lower recall, which suggests that these models are still missing some of the important rainfall signal (the physics), however, they have better precisions. The hybrid model, $\mathrm{HYB}$, improves on $\rm{NWP}$ with better recall and precision. The location-wise distribution of precision and recall across the area of interest is illustrated in Figure 6.
@@
P R F1 P R F1
CLIM 0.45 0.97 0.61 0.17 0.02 0.04
NWP 0.49 0.88 0.62 0.17 0.08 0.11
$\rm{UNET_{18}}$ 0.60 0.61 0.61 0.28 0.21 0.24
$\rm{UNET_{12}}$ 0.62 0.64 0.63 0.31 0.26 0.28
HYB 0.51 0.90 0.65 0.24 0.12 0.16
Figure 6: First row: Locationwise recall (R) of the models at threshold $\tau=0.5$mm. Second row: Locationwise precision (P) of the models at threshold $\tau=0.5$mm.
We also compare the models through their predictions of heavy rainy days, particularly on how such predictions agree with GPM-IMERG. We set the threshold for heavy rain to $\tau=10$mm. The performance of the models is much worse. This suggests that more work is needed to better predict extreme rainfall events. Figure 7 shows the location-wise distribution of P and R across the area of interest. From these figures and Table Table 4, we see again that $\rm{UNET_{12}}$ (and $\rm{UNET_{18}}$ to a lesser extent) performs better than $\rm{NWP}$. Here also the hybrid model, $\mathrm{HYB}$, has improved precision and recall compared to $\mathrm{NWP}$, but does worst than $\mathrm{UNET_{18}}$.
Figure 7: First row: locationwise precision (P) of the models at threshold $\tau=10$mm. Second row: locationwise recall (R) of the models at threshold $\tau=10$mm.